ment BACK NEXT CALCULATOR MESSAGE HY INSTRUCTOR Exercise 15-5 Bramble Inc. Issues 500 shares of $10 par value common stock and 100 shares of $100 par value preferred stock for a lump sum of $101,000 (a) Prepare the journal entry for the issuance when the market price of the common shares is $176 each and market price of the preferred is $220 each. (b) Prepare the journal entry for the issuance when only the market price of the common stock is known and it is $172 per share. $1,225. Credit account titles are automatically Indented when amount is entered. Do not indent manually. t no entry ls (Round answers to 0 decimal places, e.g. required, select No Entry for the account titles and enter O for the amounts) Debit Credit No. Account Titles and Explanation verion 4.24 hisY 疯2000-2019Jennwieyhsons.inc. Aa Rights Reserved A Division onomwetasons.is. 1 Show All
Answer #1

(Amount in $)

No. Account Title and Explanation Debit Credit
a) Cash 101,000
Common Stock 5,000
Preferred Stock 10,000
Paid In Capital - In Excess of Par Value,Common Stock 75,800
Paid In Capital - In Excess of Par Value,Preferred Stock 10,200
(Recording for Stock Issuance)
b) Cash 101,000
Common Stock 5,000
Preferred Stock 10,000
Paid In Capital - In Excess of Par Value,Common Stock 81,000
Paid In Capital - In Excess of Par Value,Preferred Stock 5,000
(Recording for Stock Issuance)

Working Notes:

a) Allocation of lump sum payment to common stock and preferred stock is shown as follows:

i) Lump sum share of common stock : (Market value of common stock / Market value of both stock) * Lump sum payment

Market value of common stock = 500 shares *$176 =$88,000

Market value of both stock = 100 shares* $220 + 500 shares *$176 = $110,000

Lump sum payment =101,000

Lump sum share of common stock: ($88,000/$110,000)*$101,000= $80,800

Paid In Capital - In Excess of Par Value,Common Stock = Lump sum share of common stock- Par value of Common Stock

=$80,800-$5,000 = $75,800

ii) Allocation of lump sum payment to preferred stock is shown as follows:

Lump sum share of common stock : (Market value of preferred stock / Market value of both stock) * Lump sum payment

Market value of preferred stock = 100 shares *$220 =$22,000

Market value of both stock = 100 shares* $220 + 500 shares *$176 = $110,000

Lump sum payment =101,000

Lump sum share of common stock: ($22,000/$110,000)*$101,000= $20,200

Paid In Capital - In Excess of Par Value,Preferred Stock = Lump sum share of Preferred stock- Par value of preferred Stock

=$20,200-$10,000 = $10,200

b)Incremental Method : As the market price of preferred stock is not known, the excess lump sum value will be distributed to it.

Common stock share in lump sum = 500 shares*$172 =$86,000

Paid In Capital - In Excess of Par Value,Common Stock= $86,000-500 shares *$10= 81,000

Paid In Capital - In Excess of Par Value,Preferred Stock = $101,000-86,000-100 shares*$100= 5,000